# A Data Generation Perspective to the Mechanism of In-Context Learning

### v2

@@ Abstract
In-Context Learning (ICL) empowers Large Language Models (LLMs) with the capacity to learn in context, achieving downstream generalization without gradient updates but with a few in-context examples. Despite the encouraging empirical success, the underlying mechanism of ICL remains unclear, and existing research offers various viewpoints of understanding. These studies propose intuition-driven and ad-hoc technical solutions for interpreting ICL, illustrating an ambiguous road map. In this paper, we leverage a data generation perspective to reinterpret recent efforts and demonstrate the potential broader usage of popular technical solutions, approaching a systematic angle. For a conceptual definition, we rigorously adopt the terms of skill learning and skill recognition. The difference between them is skill learning can learn new data generation functions from in-context data. We also provide a comprehensive study on the merits and weaknesses of different solutions, and highlight the uniformity among them given the perspective of data generation, establishing a technical foundation for future research to incorporate the strengths of different lines of research.
Machine Learning, ICML
@@ 1 Introduction
Large Language Models (LLMs) have revolutionized Natural Language Processing (NLP) (Achiam et al., 2023) and other relevant areas such as multimodality over vision and language (Liu et al., 2023a), accelerating numerous challenging research directions, e.g., AI agent (Durante et al., 2024), reasoning (Wei et al., 2022b), and story telling (Xie et al., 2023). These amazing applications are driven by LLMs’ emerging capacity of In-Context Learning (ICL), which can be formally defined as the new abilities that are not present in small models but arise in large models (Zhao et al., 2023). For instance, moral self-correction ability only occurs in LLMs of over 22B parameters (Ganguli et al., 2023).
@@
Figure 1: An illustration of ICL example on the sentiment analysis. The upper instances with the background color of gray are the labeled in-context demonstrations, while the last line is the query for which LLMs infer the sentiment label.
@@
In-Context Learning (ICL), a fundamental emerging capacity for many complicated abilities, is to utilize a few selected labeled demonstrations with the format (input, label) [^1] In this paper, we only consider classification tasks., before the query input, for making predictions in a few-/one-shot manner. An ICL example is illustrated in Figure 1.
@@
Despite the empirical success of various ICL prompting strategies for downstream applications (Mavromatis et al., 2023; Ye et al., 2022), the fundamental mechanism of ICL remains unclear, leading to unexplainable ICL phenomenons, e.g., sensitivity to the sample order (Lu et al., 2021), robust to wrong input-label mapping (Min et al., 2022). Increasing attention has been paid to understanding the ICL from various perspectives. Nonetheless, this area is still in the infant stage. Due to the complexity of LLMs, most existing works only take one individual factor into account, e.g., pre-training data distribution (Chan et al., 2022), model scale (Wei et al., 2023), and in-context task difficulties (Raventos et al., 2023). Moreover, existing works involving the same factor may adopt different experimental settings (Yoo et al., 2022; Min et al., 2022), leading to potentially conflicted conclusions. Typically, Pan (2023) details ICL as two abilities: task recognition and task learning. The task recognition ability recognizes the underlying task of demonstrations and utilizes prior knowledge embedded in weights to solve the recognized task. In contrast, the task learning ability can learn new tasks, according to ICL demonstrations, unseen in the pre-training stage. Whereas, current understanding frameworks are intuitive and ad-hoc, hindering the progress of this area. A principled and systematic view of the ICL mechanism is highly desired.
@@
In this paper, we propose the data generation perspective as a principled angle to comprehend existing research progress on the ICL mechanism. Such data generation perspective can be naturally found in both the pre-training stage and the ICL stage as the auto-regressive language model is generally utilized as the LLM backbone. Specifically, the pre-training objective is the next token prediction, aiming at modeling the data generation function of pre-training corpora; the objective of ICL inference is considered as a label generation process given the query inputs. Therefore, adopting the data generation perspective enables a unified framework through which we can cohesively address both the pre-training and the ICL stages, offering a holistic approach to understanding and leveraging the foundational mechanisms of LLMs.
Guided by the data generation perspective, we introduce a more principled and rigorous understanding framework on skill learning and skill recognition [^2] We refer to origin task learning as skill learning and origin task recognition as skill recognition, with an emphasis on the intrinsic capability of LLMs., distinguished by whether LLMs can learn a new data generation function in context. The skill learning ability learns a new data generation function in context. The new function is unseen in the pre-training stage and is (Bayesian) optimal to the given in-context demonstrations. The function learning statistical framework (Garg et al., 2022) is proposed particularly for the skill learning ability. The skill recognition ability selects one data generation function from the learned function class during pre-training. The Bayesian inference statistical framework (Xie et al., 2021) is generally utilized to understand the skill recognition ability. Notably, there exists a large gap between the two statistical frameworks, where the Bayesian inference framework utilizes the Hidden Markov Model (HMM) (Baum & Petrie, 1966) as the data generation function which is more complicated but more aligned with the real-world scenario. In contrast, the function learning framework commonly utilizes linear functions which are much simpler but distant to the real-world scenario.
Such a rigorous view also helps to ground the conceptual debate on whether the LLM can achieve intelligence into whether the LLM can learn new data generation functions in context, clearly unveiling the strengths and limitations of current LLM approaches to intelligent systems.
Our other contribution is, given the data generation perspective, we shed light on the transferability of two separated research lines of Bayesian inference and function learning. Specifically, we extend the Bayesian inference framework as a technical solution for skill learning, therefore opening the door to incorporate real-world cases, e.g., topics of corpora.
Organization: We introduce skill recognition with the Bayesian inference framework and skill learning with the function learning framework in Section 2 and 3, respectively. We then showcase the uniformity of the two frameworks via extending the original Bayesian inference framework for skill learning in Section 4. More empirical analysis on skill learning and skill recognition abilities are illustrated in Section 5. Future directions are discussed in Section 6.
@@
Comparison with existing relevant literature. As far as we know, this paper is the first to provide a comprehensive discussion on existing studies about the mechanism of ICL and advocating a principled data generation perspective. This paper is different from existing surveys such as Dong et al. (2022); Zhao et al. (2023); Wei et al. (2022a).
Terminology. The prompt sequence of In-Context Learning consists of two parts: (1) The demonstrations illustrate as (input, label) pair, denoted as $(x_{i},y_{i})$; The demonstrations provide the basic description of the intended task. (2) The query is the test input after a few demonstrations. The ICL aims to provide the correct prediction for the query based on the in-context demonstrations and the knowledge of pre-trained LLM. The data generation function in this paper refers to the underlying hypothesis on textual data generation. It serves as the data assumptions in the theoretical understanding and the simulation data generator for the synthetic experimental analysis. Each data generation function obtained by the LLM can be recognized as a skill.
@@ 2 Skill Recognition
The skill recognition ability is to select the most proper data generation function from the function class obtained during pre-training, according to in-context demonstrations. A Bayesian inference framework (Xie et al., 2021) is introduced to explain skill recognition. The ICL inference can then be instantiated as a Bayesian inference process as follows:
$\small\begin{split}&p(\text{output}|\text{prompt})=\\ &\int_{\text{concept}}p(\text{output}|\text{concept},\text{prompt})p(\text{concept}|\text{prompt})d(\text{concept})\end{split}$ (1)
@@
where $p(\text{output}|\text{prompt})$ is the conditional probability of the output generation given the prompt. It can be marginalized with the pre-training concepts and each concept corresponds to a pre-training data generation function. $p(\text{concept}|\text{prompt})$ is to locate the latent concept aligned with the in-context demonstrations, which will be the key focus in the ICL inference stage. After locating the aligned concept, $p(\text{output}|\text{concept},\text{prompt})$ is to utilize the selected data generate function for the output generation.
@@
Such modeling method of latent concept is commonly adopted in the NLP domain as the language data is inherently compositional with unobserved latent concepts that are not explicitly present in the raw text, e.g., sentiments, topics, and syntactic structures (Chung et al., 2015; Zhou et al., 2020). Latent variable models can specify prior knowledge and structural dependency for language data which enjoys the characteristics of high compositionality. Deep latent variable models are popularly utilized to improve various tasks such as the alignment in statistical machine translation, topic modeling, and text generation (Kim et al., 2018; Fang et al., 2019; Wang et al., 2023).
@@
However, there are various definitions of the aforementioned concept, all latent information that can help ICL to make the inference could be considered as a good choice of concept. We summarize the existing concept definitions as follows: (1) Xie et al. (2021) defines the concept as the transition matrix $\mathbf{\theta}$ of a Hidden Markov Model (HMM) (Baum & Petrie, 1966). The concept helps to state a transition distribution over observed tokens. A concrete example of the concept is the transition between name (Albert Einstein) $\to$ nationality (German) $\to$ occupation (physicist) ) in wiki bios. (2) Wang et al. (2023) simplifies the transition between tokens, modeled by HMM, with LDA topic models where each topic corresponds to one latent concept (Blei et al., 2003). However, the LDA topic model cannot capture the dependency between tokens. (3) Despite the above mathematical interpretations, Todd et al. (2023); Liu et al. (2023b) empirically establish the connection between the latent concept and the downstream task, e.g., supervised classification and question-answering, where the particular latent representation in the LLM can capture essential information about the task. The following discussions focus on the first concept definition (Xie et al., 2021) as the transition matrix of HMM.
@@
Xie et al. (2021) utilizes the Bayesian inference framework to understand how the obtained pre-training data functions are activated by the in-context demonstrations. The key questions of this framework are: (1) In the pre-training stage, how the model obtained the latent concepts from the pre-training corpus; and (2) In the ICL inference stage, how the in-context demonstrations can locate the most relevant concept to generate the desired output for the query.
@@
During the pre-training stage, the LLM can obtain various concepts from the large pre-training corpora if each pre-training document is generated from an individual HMM model. In such case, the next token prediction objective can converge if and only if the LLM can successfully generate the correct next token matching the HMM generation process dominated by the underlying concept. Notably, different documents can be generated from various concepts sampled from the concept set $\Theta$.
@@
The ICL inference stage conducts an implicit Bayesian inference to locate the aligned pre-training concept $\theta^{*}\in\Theta$ with the optimal likelihood to generate the observed in-context demonstrations. The format of the prompt is shown below:
@@
$\small\left[S_{n},x_{\text{test}}\right]=\left[x_{1},y_{1},o^{\text{del}},\ldots,x_{n},y_{n},o^{\text{del}},x_{\text{test}}\right]\sim p_{\text{prompt}}$ (2)
@@
where $p_{\text{prompt}}$ is a data generation process implemented with HMM parameterized by the $\theta^{*}$. $x_{i}$, $y_{i}$ and $o^{\text{del}}$ are the input, label, and delimiter, respectively. The difficulty in locating $\theta^{*}$ is that it is of low probability for all the concepts to generate the in-context demonstrations. The key reason is that token transition patterns of the in-context demonstrations are in three types (1) the input to the label $x_{i}\to y_{i}$ (2) the label to the delimiter (3) the delimiter to the input. The latter two patterns hardly appear in the pre-training concepts due to different delimiter usage.
@@
To address the above issue of low probability, the authors propose some strong assumptions that may not be feasible in real-world scenarios. We showcase one assumption as follows. It is more likely for the selected concept $\theta^{*}$ to transit to the delimiter hidden states while other concepts are unlikely to transit to the delimiter hidden states. Equipped with those assumptions, we are able to locate the aligned pre-training concept to implement Bayesian inference. Even though we cannot locate the aligned concept, Xie et al. (2021) provides the theoretical guarantee on the effectiveness of the ICL in such cases, where the ICL performance improves along with the in-context example length.
@@
Inspired by the above Bayesian inference framework, more frameworks on understanding skill recognition are proposed, e.g., the PAC-Bayesian framework and Hopfield Network. Zhang et al. (2023b) further takes the architecture of transformers into consideration and analogies ICL inference to a Bayesian model averaging algorithm. Wies et al. (2023) propose a PAC-based generalization framework exhibiting satisfying generalization bound on the ICL where a transformer trained on multi-task can match the ICL performance of a transformer trained solely on the downstream task. Zhao (2023) analogizes the latent concept location as memory retrieval with the Hopfield Network. Nonetheless, all the existing studies are either theoretical or empirical based on synthetic data without considering the real-world scenario. It could be a promising future direction to investigate how the LLM is able to retrieve the concept with natural language explanations.

### v3

# A Survey to Recent Progress Towards Understanding In-Context Learning

@@ Abstract
In-Context Learning (ICL) empowers Large Language Models (LLMs) with the ability to learn from a few examples provided in the prompt, enabling downstream generalization without the requirement for gradient updates. Despite encouragingly empirical success, the underlying mechanism of ICL remains unclear. Existing research remains ambiguous with various viewpoints, utilizing intuition-driven and ad-hoc technical solutions to interpret ICL. In this paper, we leverage a data generation perspective to reinterpret recent efforts from a systematic angle, demonstrating the potential broader usage of these popular technical solutions. For a conceptual definition, we rigorously adopt the terms of skill recognition and skill learning. Skill recognition selects one learned data generation function previously seen during pre-training while skill learning can learn new data generation functions from in-context data. Furthermore, we provide insights into the strengths and weaknesses of both abilities, emphasizing their commonalities through the perspective of data generation. This analysis suggests potential directions for future research.
A Survey to Recent Progress Towards Understanding In-Context Learning
Haitao Mao1, Guangliang Liu1, Yao Ma2, Rongrong Wang1, Kristen Johnson1, Jiliang Tang1 1Michigan State University 2Rensselaer Polytechnic Institute {haitaoma,liuguan5,wangron6,kristenj,tangjili}@msu.edu may13@rpi.edu
@@ 1 Introduction
LLMs have revolutionized Natural Language Processing (NLP) Achiam et al. (2023) and other relevant areas such as multi-modal tasks over vision and language Liu et al. (2023a), accelerating numerous challenging research directions, e.g., AI agent Durante et al. (2024), reasoning Wei et al. (2022b), and story telling Xie et al. (2023). These amazing applications display LLMs’ emerging capabilities, which can be formally defined as new abilities that are not present in small models but arise in larger ones Zhao et al. (2023). Among them, the emerging ICL ability serves as an important foundation of other capabilities. Notably, small models also have the capability to perform ICL, but the level of capability is different from that of larger models, wherein people can easily observe more in-depth displays of understanding for the given context of inputs, e.g., identify long-term dependency and abstract concept comprehension. For instance, Ganguli et al. (2023) demonstrates that only LLMs over 22B parameters can understand the moral concepts, being able to generate unbiased answers.
@@
Figure 1: Illustration of ICL for Sentiment Analysis. The upper instances (with background color gray) are the labeled in-context demonstrations, while the last line is the query for which LLMs infer the sentiment label.
@@
ICL, a fundamental and emerging capability serving as the pre-requisite for many complicated abilities, is the process of leveraging a few selected labeled demonstrations with the format (input, label) [^1] In this paper, we focus on classification tasks, as they are widely used in theoretical studies of ICL due to their well-defined mathematical tools and clear evaluation metrics., before the query input, for making predictions in a few-/one-shot manner. An example of ICL is illustrated in Figure 1.
Table 1: A summarization table of representative works. SR and SL stand for skill recognition and skill learning, respectively. Function approximation revolves on how effectively ICL can fit different generalize functions. The Internal Mechanism describes how LLMs learn through various gradient descent algorithms. More details on empirical simplification and theoritical assumptions can be found in Appendix D.
Literature Ability Analysis View Date Generation Function Characteristics
Xie et al. (2021); Zhang et al. (2023c) SR Theoretical HMM Internal Mechanism
& Empirical
Wang et al. (2023) SR Empirical LDA Generalization
Zhao (2023) SR Theoretical Hopfield Network Internal Mechanism
Raventos et al. (2023) SL Theoretical linear regression Generalization
Wu et al. (2023a) SL Empirical linear regression Generalization
Garg et al. (2022) SL Empirical linear regression, decision tree, NN Function Approximation
Bai et al. (2023); Fu et al. (2023a) SL Theoretical linear regression, decision tree, NN Generalization
Yadlowsky et al. (2023); Ahuja et al. (2023) SL Empirical linear regression, polynomial regression Generalization
Von Oswald et al. (2023); Zhang et al. (2023b) SL Theoretical linear regression Internal Mechanism
Mahankali et al. (2023); Ahn et al. (2023a)
Akyürek et al. (2022) SL Theoretical linear regression Internal Mechanism
Li et al. (2023a); Ren and Liu (2023) SL Theoretical non-linear regression Internal Mechanism
Cheng et al. (2023); Guo et al. (2023)
Hahn and Goyal (2023) SR&SL Theoretical context-free grammar Generalization
@@
Despite the empirical success of various ICL prompting strategies for downstream applications Mavromatis et al. (2023); Ye et al. (2022), the mechanism of ICL remains unclear, leading to unexplainable observations, e.g., sensitivity to the sample order Lu et al. (2021), or being robust to human-crafted yet irrational input-label mapping. Increasing attention has been paid to understand ICL from various perspectives. However, this area is still growing, with many open research questions are actively being explored. Due to the complexity of LLMs, most existing works only take one individual factor into account, e.g., the pre-training data distribution Chan et al. (2022a), model scale Wei et al. (2023), or difficulty level of the in-context task Raventos et al. (2023). Moreover, existing works focusing the same factor may adopt different experimental settings Yoo et al. (2022); Min et al. (2022), leading to potentially conflicting conclusions. Typically, Pan (2023) categorizes ICL into two abilities: task recognition and task learning.
@@
In this paper, we propose the data generation perspective as a principled angle to comprehend existing studies towards understanding ICL. Following this perspective, the pretraining stage can be interpreted as learning the data generation function classes underlying pretraining corpus, where the masked language modeling objective Devlin et al. (2019) and the next token prediction objective Radford et al. (2018) are both objectives that allow us learn the data generation functions. Similarly, the ICL stage can be considered as a label generation process given the query inputs. Therefore, adopting this data generation perspective enables a unified framework through which we can cohesively analyze both pretraining and ICL stages, offering a holistic approach to understanding the foundations of LLMs.
Guided by the data generation perspective, we introduce a more principled and rigorous understanding framework on skill learning and skill recognition, distinguished by whether LLMs can learn a new data generation function in context. The skill learning ability is to learn a new data generation function in context, which is unseen in the pretraining stage. The skill recognition ability selects one learned data generation function previously seen during pre-training. To analyze the mechanism of abilities, the function learning statistical framework Garg et al. (2022) and the Bayesian inference statistical framework Xie et al. (2021) are representative works for skill learning and skill recognition ability, respectively.
Organization: Section 2 introduces previous studies of ICL and Section 3 presents the terminology. Key contributions lie in Section 4 and 5, which systematically review the skill recognition with the Bayesian inference framework and the skill learning with the function learning framework, respectively. We outline the challenges and potential directions in Section 6, aiming to offer a valuable guide for newcomers to the field while also illuminating pathways for future research.
2 Related works
Comparison with existing relevant literature. The key difference between our work and existing ones lies in its more dedicating scope on the mechanism of ICL and advocating a principled data generation perspective, instead of a broad, application-oriented perspective in Dong et al. (2022); Zhao et al. (2023); Wei et al. (2022a). Our work provides a comprehensive literature review and clear categorization as shown in Table 1. Moreover, we propose a new holistic data generation perspective which can be utilized for the tokenizer, which clarifies the connections and distinctions between different data statistical frameworks.
@@ 2 Related works
Distinguish skill recognition/learning from task recognition/learning Pan (2023). We distinguish our proposed skill recognition/learning from a data generation perspective with previous task recognition/learning proposed in Pan (2023). Task recognition/learning is a narrower aspect of our skill recognition/learning as they majorly focus on the empirical performance variation under the label permutation on in-context data. Task learning is recognized as performance degradation, indicating ICL learns the permuted in-context data. In contrast, the task recognition corresponds to the unchanged performance, indicating ICL only relies on pre-training knowledge. The key advantages of our proposed skill recognition/learning definition are shown as follows: (1) Thanks to the mathematical description with a data generation function, skill learning/recognition enables both theoretical analysis and empirical evidence, instead of only focusing on the empirical one. (2) Task recognition/learning can only emphasize the performance of a classification task in complicated real-world applications. Instead, skill learning/recognition can utilize different existing data generation functions in the NLP domain, e.g., HMM, and LDA, rather than merely input-label mapping for classification. Moreover, the data generation enables to conduct synthetic analyses in a systematic and controllable setting.
3 Terminology
The prompt sequence of In-Context Learning consists of two parts: (1) The demonstration is illustrated as an (input, label) pair, denoted as $(x_{i},y_{i})$; These demonstrations provide the basic description of the intended task. (2) The query is the test input after a few demonstrations. ICL aims to provide the correct prediction for the query based on the in-context demonstrations and the prior knowledge of a pre-trained LLM. The data generation function in this paper refers to the underlying hypothesis on language data generation. It serves as the data assumption in the theoretical understanding and the simulation data generator for the synthetic experimental analysis. Each data generation function obtained by the LLM can be recognized as a skill.
@@ 4 Skill Recognition
Skill recognition ability is the ability of an LLM to select the most proper data generation function from the function class obtained during pre-training. And this selection process is driven by the in-context demonstrations. A Bayesian inference framework Xie et al. (2021) is introduced to explain the skill recognition. The ICL inference can be instantiated as a Bayesian inference process as follows:
$\begin{split}&p(\text{y}|\text{prompt})=\\ &\int_{\text{concept}}\!\!p(\text{y}|\text{concept},\!\text{prompt})p(\text{concept}|\text{prompt})d(\text{concept})\end{split}$
@@
where $p(\text{y}|\text{prompt})$ is the conditional probability of the output generation y given the prompt. It can be marginalized with pre-training concepts and each concept corresponds to a pre-training data generation function. $p(\text{concept}|\text{prompt})$ is the probability of locating the latent concept aligned with in-context demonstrations. After locating the aligned concept, $p(\text{y}|\text{concept},\text{prompt})$ utilizes the selected data generation function for the output generation.
@@
This approach to modeling latent concepts is widely used in the field of NLP, as language data is inherently compositional, involving underlying concepts—such as sentiment, topics, and syntactic structures—that are not explicitly observable in the raw text Chung et al. (2015); Zhou et al. (2020). Latent variable models can specify prior knowledge and structural dependencies for language data which enjoys the characteristics of high compositionality. Deep latent variable models are popularly utilized to improve various tasks such as alignment in statistical machine translation, topic modeling, and text generation Kim et al. (2018); Fang et al. (2019); Wang et al. (2023).
@@
Though there are various definitions of latent concepts, any latent information that can help ICL can be considered as a good choice for the concept in the Bayesian inference process above. We summarize the existing concept definitions as follows: (1) Xie et al. (2021) defines the concept as the transition matrix $\mathbf{\theta}$ of a Hidden Markov Model (HMM) Baum and Petrie (1966), which assumes to be the underlying distribution of the real-world language data. The concept helps to state a transition distribution over observed tokens. A concrete example of the concept is the transition between name (Albert Einstein) $\to$ nationality (German) $\to$ occupation (physicist) in wiki bios. (2) Wang et al. (2023) simplifies the transition between tokens, modeled by HMM, with LDA topic models where each topic corresponds to one latent concept Blei et al. (2003). (3) Despite the above mathematical interpretations, Todd et al. (2023) and Liu et al. (2023b) empirically establish the connection between the latent concept and the downstream task, e.g., supervised classification and question-answering, where the particular latent representation in the LLM can capture essential information about the task.
@@
The Bayesian inference framework is firstly proposed by Xie et al. (2021), interpreting how obtained pre-training data functions are activated by in-context demonstrations. Key challenges in this framework are: (1) In the pre-training stage, how the model obtains the latent concepts from the pre-training corpus; and (2) In the ICL inference stage, how in-context demonstrations can locate the most relevant concept to generate the desired output.
@@
The pre-training stage aims to obtain various concepts from the large pre-training corpora if each pre-training document is generated from an individual HMM model. In such cases, the next token prediction objective can converge if and only if the LLM can successfully generate the correct next token matching the HMM transitions. The transitions are dominated by the underlying concept Xie et al. (2021). Different documents can be generated from various concepts sampled from the concept set denoted as $\Theta$.
@@
The ICL inference stage conducts an implicit Bayesian inference to locate an appropriate concept $\theta^{*}\in\Theta$ which shows the optimal likelihood to generate the given in-context demonstrations. The format of the prompt is shown below:
@@
$\begin{split}&\left[S_{n},x_{\text{test}}\right]\\ &=\left[x_{1},y_{1},o^{\text{del}},\ldots,x_{n},y_{n},o^{\text{del}},x_{\text{test}}\right]\sim p_{\text{prompt}}\end{split}$ (1)
@@
where $p_{\text{prompt}}$ is a data generation process implemented with HMM parameterized by $\theta^{*}$. $x_{i}$, $y_{i}$ and $o^{\text{del}}$ are the input, label, and delimiter, respectively. The difficulty in locating $\theta^{*}$ is due to low probability for all the pre-training concepts to generate the in-context demonstrations. The key reason is that token transition patterns of the in-context demonstrations are of three types: (1) the input to the label $x_{i}\to y_{i}$, (2) the label to the delimiter, and (3) the delimiter to the input. The latter two patterns hardly appear in the pre-training data due to different delimiter usages.
@@
To address the above issue of low probability, Xie et al. (2021) proposes some assumptions. One example is the located concept $\theta^{*}$ enjoys a higher probability transiting to delimiters than that of other concepts. Equipped with those assumptions, we are able to locate the aligned pre-training concept to implement Bayesian inference. The model can locate the corrrect concept with $p(\theta^{*}|\text{prompt})=1$ and $p(\theta|\text{prompt})=0$ for all $\theta\in\Theta\setminus\theta*$. Even though we cannot locate the aligned concept, Xie et al. (2021) provides the theoretical guarantee on the effectiveness of the ICL in such cases, where the ICL performance improves along with the increasing number of in-context examples.
@@
Inspired by the above Bayesian inference framework, more methods towards understanding skill recognition are proposed, e.g., the PAC-Bayesian framework Alquier et al. (2024) and Hopfield Network Hopfield (2007). Zhang et al. (2023c) analogizes ICL inference to a Bayesian model averaging algorithm. Wies et al. (2023) presents a PAC-based generalization framework exhibiting satisfying generalization bound on the ICL where a transformer trained on multi-task can match the ICL performance of a transformer trained solely on the downstream task. Zhao (2023) analogizes the latent concept location as memory retrieval with the Hopfield Network. More recently, a novel information-theoretic framework Jeon et al. (2024) has been introduced, decomposing the ICL prediction error into three distinct terms: irreducible error, meta-learning error, and intra-task error. This decomposition helps aligning ICL with existing studies hypothesizing ICL as an instance of meta-learning.
Nonetheless, existing studies are based on either synthetic data or pure theoretical analysis. It could be a promising direction to investigate how LLMs retrieve concepts and how to interpret the retrieved concept through natural language.